# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v1

@@ 4.4. Practicality with Consumer-Grade Channels
Despite advancements in brainwave sensors in recent years, consumer-grade devices still have a limited number of sensors. Therefore, we selected three devices to evaluate performance by focusing exclusively on the electrodes specific to each device. The selected devices include the Emotiv Epoc X, which features 14888AF3, F7, F3, FC5, T7, P7, O1, O2, P8, T8, FC6, F4, F8, AF4 electrodes and is priced at 999 dollars, and the Muse 2, equipped with 4999TP9, AF7, AF8, TP10 electrodes and available at 269.99 euros, offer a budget-friendly options for consumer use. Additionally, we considered the DSI-VR300101010FCz, Pz, P3, P4, PO7, PO8, Oz, which includes 7 electrodes and costs approximately 10,000 dollars. This device, though expensive, has the potential to become more affordable in the future and could be integrated into XR devices.
@@
We selected the same electrodes as used in the devices and trained a model for each device. Table 7 represents the results with a lower number of channels. The baseline EER increases from 10.78% to 13.97%, 16.55%, and 21.94% for Emotiv Epoc X, DSI-VR300, and Muse 2, respectively. However, considering the 93 channels in our main dataset and reducing them to 14, 7, and 4 channels indicates some redundant information in the original channels due to EEG correlations across electrode recordings. Moreover, using averaging techniques and multiple enrollment sessions can further improve the results.
Table 8. Summary of state-of-the-art brain biometric verification solutions:
Publication Subjects Sess. Sess. Interval Verification time public dataset open source EER (%)
Das et al.  (Das et al., 2016), 2016 50 150 up to 49 days 30s $\times$ $\times$ 10
Wu et al.  (Wu et al., 2018), 2018 40 80 30 days 6s $\times$ $\times$ $5.75<\text{EER}<7.07$
Maiorana et al.  (Maiorana and Campisi, 2017), 2018 45 270 up to 36 months 5s $\times$ $\times$ $6.6\leq\text{EER}\leq 10.7$
Maiorana (Maiorana, 2021), 2021 45 225 up to 15 months 5s $\times$ $\times$ $8\leq\text{EER}\leq 10.7$
Debie et al. (Debie et al., 2021), 2021 54 108 different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. $1s-4s$$\times$ $1.32<\text{EER}<2.22$
Chaurasia et al. (Chaurasia et al., 2024), 2024 54 108 different days [^3] The paper did not specify the exact number of days, so we used the same terminology. However, it typically refers to a very short time interval. 1s 11.99
Proposed Method 345 6013 up to 66 months 1s 10.78
Proposed Method 345 6013 up to 66 months 4s 7.98
4.5. Comparative Analysis
@@
We first evaluate the proposed method against state-of-the-art brainwave authentication approaches. Subsequently, we analyze its performance relative to international biometric standards to provide context and clarify the significance of the results.
4.5.1. Compare with Brainwave Authentications Systems
Table 8 summarizes the state-of-the-art in multi-session brainwave authentication studies. As shown, previous studies typically involve 40–54 subjects and 80–270 sessions, which is significantly lower than our dataset of 345 subjects and 6013 sessions. From the perspective of time intervals between sessions, prior works vary: some datasets span a few days, such as those in (Chaurasia et al., 2024; Debie et al., 2021), which report inconsistent EER values ranging from 1.32% to 11.99%. Others span 1–2 months with EER values between 5.75% and 10%(Das et al., 2016; Wu et al., 2018), and the study by Maiorana et al.(Maiorana, 2021; Maiorana and Campisi, 2017) includes intervals of up to 36 months with EER values between 6.6% and 10.7%. However, our dataset includes time intervals of up to 66 months with EER values of 7.98% and 10.78%.
While we have already demonstrated that a lower number of subjects or sessions does not necessarily make authentication evaluation easier, we observe that evaluation with a lower number of subjects could have higher variability, making it more susceptible to overfitting. We also observed that longer time intervals between sessions can significantly increase the error rate. Therefore, in terms of results, we can conclude that our findings are comparable to or outperform the state-of-the-art, despite using a more challenging dataset and shorter verification times. However, only Debie et al. (Debie et al., 2021) show a lower EER compared to our result, even when comparing our result after a day (7.7%), where they train a deep model network per subject based on a single session. While this approach is not scalable as it requires training a CNN per subject, it is not obvious how the model could learn session variants while learning only from a single session and evaluated by the second one, and the source code was not public. But Chaurasia et al.(Chaurasia et al., 2024) used the same dataset and reported 20.82% EER based on triplet loss function and Euclidean distance, and 11.99% based on PSD as feature extraction and SVM as feature compression, which shows the data was not enough to take advantage of deep learning even based on multiple subjects. Additionally, our study is the only one to use a publicly available dataset with time intervals exceeding a few days. It is also the only one, except Chaurasia et al.(Chaurasia et al., 2024), to release its source code on multi-session brainwave authentication.
@@ 4.5.1. Compare with Brainwave Authentications Systems
Table 9. Performance metrics for different scenarios with detailed enrollment and verification session information.
@@
First All Others 1 $10.78\pm 6.67$ $52.82\pm 24.54$ $79.77\pm 18.41$ $91.97\pm 11.43$
First All Others 4 $7.98\pm 7.00$ $40.97\pm 29.46$ $67.56\pm 26.29$ $81.86\pm 21.59$
First Second 1 $7.67\pm 7.28$ $45.77\pm 33.58$ $73.51\pm 29.72$ $84.76\pm 23.37$
First Second 4 $4.23\pm 6.39$ $35.10\pm 38.63$ $56.06\pm 41.13$ $66.15\pm 38.80$
First Two Third 1 $6.01\pm 5.29$ $35.98\pm 28.36$ $66.49\pm 27.72$ $81.57\pm 22.49$
First Two Third 4 $2.67\pm 3.87$ $22.67\pm 33.28$ $44.78\pm 38.92$ $57.59\pm 38.31$
First Two Third 16 $1.41\pm 2.70$ $21.13\pm 37.59$ $31.87\pm 41.42$ $38.96\pm 42.80$
4.5.2. Comparison with Biometric Industry Standards
Brainwave authentication is still in its early stages, making it crucial to assess its current progress and identify the gaps for real-world implementation. For a biometric system to be deployed in practical applications, it must adhere to international industrial biometric standards. These standards emphasize a very low FAR to ensure security, while maintaining a reasonable FRR to keep the system usable for legitimate users. According to guidelines from NIST (2023)/ISO [^11] https://pages.nist.gov/800-63-3/sp800-63b.htmland the European Border Guard Agency Frontex [^12] https://www.frontex.europa.eu/assets/Publications/Research/Best_Practice_Technical_Guidelines_ABC.pdf, biometric systems are required to operate at FAR $\leq 0.1\%$ (1 in 1,000). Stricter recommendations from FIDO [^13] https://fidoalliance.org/specs/biometric/requirements/Biometrics-Requirements-v4.0.1-fd-20240522.pdf and the upcoming NIST (August 2024)/ISO 2024 standards [^14] https://pages.nist.gov/800-63-4/sp800-63b.html suggest a FAR $\leq 0.01\%$ (1 in 10,000). For usability, FRR standards propose a maximum FRR of $\leq 5\%$, allowing at least 19 successful logins out of 20 attempts for genuine users.
Table 9 presents the FRR at low False FAR of 1%, 0.1%, and 0.01%. The results show higher FRR. For instance, with two enrollment sessions and a verification time of 4 seconds, the FRR is 44.78% at FAR 0.1% and 57.59% at FAR 0.01%, meaning approximately 11 and 8 successful logins out of 20 attempts for legitimate users, respectively. However, we also observed very high standard deviations, indicating that the system works well for some subjects but performs worse for others. For example, under the same scenario, the FRR is less than or equal to 20% for 41 and 26 subjects at FAR 0.1% and is less than or equal to 5% for 30 and 14 subjects at FAR 0.01%. These observations highlight the need for further investigation at the subject level.
@@ 4.5.2. Comparison with Biometric Industry Standards
While our experiments show that brainwave authentication does not yet meet international standards, our previous experiment on the impact of headsets on performance (Sec.4.2.4) revealed that the headset with a higher number of subjects yielded a lower error rate, and we also know that biometrics that meet these standards typically employ datasets with thousands of subjects(Guo et al., 2016). Therefore, we conducted an experiment to explore the relationship between the number of subjects used for training the feature extractor and the error rate. The results, presented in Figure 8, indicate a linear relationship between the logarithm of the number of subjects and the EER. Based on this relationship, we predict that if 1,024–2,048 subjects were available for training, the EER could drop below 0.5%, which would satisfy the required standards (Fallahi et al., 2024a).
@@
Figure 8. Relationship between the number of subjects used to train the feature extractor and the EER. The plot shows experimental results (blue points with error bars) and a linear prediction (dashed line) based on these results.
5. Limitations
@@
Our study on the effective parameters influencing EEG-based authentication performance in a multi-session scenario is subject to certain limitations, discussed below. First, our stimuli were limited to a general text-based memorization task. Ideally, multiple task types should have been included to allow for performance comparisons and identification of the best-performing task. Second, we simulated consumer-grade devices by selecting the same number of channels in the same locations as those found in such devices. However, we did not have a measure to directly compare the quality of sensors. Additionally, we did not consider downsampling the data to simulate sensor quality, as downsampling served as a noise reduction method in multi-session authentication (Maiorana, 2021). It is worth noting that all other multi-session studies in the literature (Maiorana and Campisi, 2017; Debie et al., 2021; Wu et al., 2018; Chaurasia et al., 2024) also employ medical-grade data. Finally, we had access to only a single dataset that included a notable number of sessions and subjects. However, best practices in state-of-the-art face recognition research recommend training on multiple large datasets and testing on an independent dataset to ensure a more robust evaluation (Kim et al., 2022). Achieving this requires massive amounts of data to extract distinctive features while accounting for subject, environmental, and hardware-specific effects. For example, Kim et al. (Kim et al., 2022) used three datasets comprising over 350,000 subjects and more than 14.9 million images to train a generalized model capable of evaluation on a few independent datasets.
6. Conclusion and Future Work
@@
This study aimed to investigate the factors influencing the performance of brainwave authentication over multiple sessions. We conducted evaluations using the largest multi-session brainwave authentication dataset available to date, which includes 345 subjects and 6,013 sessions. Our findings highlight that a pairwise deep learning approach for feature extraction combined with cosine similarity for feature comparison is the most effective method for processing brainwave data. Interestingly, we observed that larger test sets do not necessarily make authentication more difficult; instead, they reduce the uncertainty of the results. Authentication performance remained relatively stable during the first week, with error rates increasing only slightly in the second month. However, error rates could increase by up to 2.5 times after one year, suggesting that the enrollment dataset should be periodically updated to maintain performance. We also examined the impact of recording hardware on results, emphasizing the importance of including data from different headsets in training the feature extractor. Additionally, we investigated the trade-offs affecting performance and showed that increasing the number of enrollment sessions and verification time can help reduce the EER, though only up to a certain point and at the cost of time. Our findings also demonstrate that brainwave authentication remains feasible even with a reduced number of EEG sensors. Finally, we demonstrate that our results are competitive with the state-of-the-art in brainwave authentication, investigated with stronger assumptions in simpler scenarios. However, achieving industrial standard-level performance would require collecting data from a larger number of subjects.
@@
For future work, there is an urgent need for a public multi-session dataset with more than 1,024 subjects to advance research in brainwave authentication. This dataset should be collected using consumer-grade devices in multiple sessions, and incorporating multiple devices in the same session would be valuable for studying the cross-headset authentication challenge. On the methodological side, further research should focus on developing deep learning architectures that better capture the unique features of brainwave data, as well as effective signal preprocessing approaches to remove noise. Adopting an open-source approach would enable researchers to build on each other’s work, promoting reproducibility and collaboration rather than isolated, non-reproducible studies.
Acknowledgements.
@@
This work was funded by the Topic Engineering Secure Systems of the Helmholtz Association (HGF) and supported by KASTEL Security Research Labs, Karlsruhe, and Germany’s Excellence Strategy (EXC 2050/1 ‘CeTI’; ID 390696704). Portions of the research in this paper use the FERET database of facial images collected under the FERET program.

### v2

@@ IV-D Practicality with Consumer-Grade Channels
Despite advancements in brainwave sensors in recent years, consumer-grade devices still have a limited number of sensors. Therefore, we selected three devices to evaluate performance by focusing exclusively on the electrodes specific to each device. The selected devices include the Emotiv Epoc X, which features 14888AF3, F7, F3, FC5, T7, P7, O1, O2, P8, T8, FC6, F4, F8, AF4 electrodes and costs 999 dollars, and the Muse 2, equipped with 4999TP9, AF7, AF8, TP10 electrodes and available at 269.99 euros. Both are budget-friendly options for consumer use. Additionally, we considered the DSI-VR300101010FCz, Pz, P3, P4, PO7, PO8, Oz, which includes 7 electrodes and costs 10,000 dollars. This device, though expensive, has the potential to become more affordable in the future and could be integrated into XR devices.
@@
We selected the same electrodes as used in the devices and trained a model for each device. Table IX represents the results with a lower number of channels. The baseline EER increases from 10.28% to 13.37%, 16.61%, and 18.18% for Emotiv Epoc X, DSI-VR300, and Muse 2, respectively. The reduction from 93 channels in our main dataset to 14, 7, and 4 channels suggests redundancy due to correlated EEG signals across electrodes. Moreover, using averaging techniques and multiple enrollment sessions can further improve the results.
IV-E Comparative Analysis
@@
We first evaluate the proposed method against state-of-the-art brainwave authentication approaches. Next, we analyze its performance relative to international biometric standards to provide context and clarify the significance of the results.
IV-E1 Comparison with Brainwave Authentications Systems
Table VI summarizes the state-of-the-art in multi-session brainwave authentication studies.
Das et al. [10] is one of the earliest multi-session studies, where the authors used simple averaging of samples and compared them using cosine distance. Their reported EER ranges between approximately 10% and 22%. The best result for a one-week interval is around 10% EER, while for an average interval of 34 days, the EER is approximately 15%.
Wu et al.[11] and Debie et al.[13] do not report EER directly. Wu et al. report an FRR of 7.07% and FAR of 5.75% with a 6-second window, while Debie et al. report an FRR of 1.93% and FAR of 1.16% using 7.3 seconds of data with only a few days between sessions. Debie et al.[13] performance is calculated by averaging two scenarios: (1) first session as enrollment and second as verification, and (2) the reverse—an unrealistic setup that assumes access to future data. Both studies use two-session datasets and train a separate CNN per subject, making it unclear how session variability is learned. This design limits generalization and requires training a new model for each user, which is impractical and does not support inter-subject learning.
Maiorana et al. [16, 12] present a longitudinal evaluation with 45 subjects over six sessions collected across three years. In their first study, they used handcrafted features with a Hidden Markov Model and reported EERs between 6.6% and 10.7% for 5-second verification samples. In the second study, the evaluation was limited to 15 months without a clear explanation. They adopted a metric learning approach, training a feature extractor using contrastive loss on 30 subjects, with a separate CNN per channel, and used a one-class SVM for classification. They reported EERs between 4.8% and 10.7% based on 15 evaluation subjects. However, this approach has key limitations: learning from single channels ignores inter-channel relationships, and one-class SVMs have shown limited reliability, as supported by our findings.
Chaurasia et al.[14] present a framework for EEG evaluation that investigates the multi-session scenario. They show that, on a small dataset, handcrafted features outperform deep learning approaches , reporting an EER of 11.99% using PSD and AR features combined with an SVM, compared to 20.68% EER using a CNN with triplet loss.
In terms of results, when using the first session for enrollment and the second for verification—with an average interval of 28 days—we achieve an EER of 3.04%, outperforming [10, 11, 14]. Maiorana et al.[16, 12] report 4.8% EER with a 5-second window and similar time interval (a month), while we obtain 3.04% with a 4-second window. Compared to Debie et al. [13], who report EER between 1.93% and 1.16% using a 7.3-second window and only a few days between sessions, our 2.30% EER with a 7-second window is measured across a average of 28-day interval. Overall, our results either outperform or are competitive with the state of the art.
However, several important issues are overlooked in related works: (1) Most of the studies use short intervals [10, 11, 13, 11] between sessions while our work considers longer intervals, making the evaluation more realistic for practical applications. (2) Except for Chaurasia et al.[14], none of the studies provide source code, making it impossible to reproduce results or build upon them. This lack of transparency makes it particularly difficult to verify findings, especially where reported performance differs from our results and expectations, such as the use of one-class SVM [12] and subject-specific CNNs [13, 11] (3) None of the reviewed studies examine the effect of recording hardware on authentication performance, which is a key factor for real-world deployment. (4) Some related works [10, 11, 13, 16] do not follow an inter-subject approach, where a general feature extractor is trained to avoid the need for handcrafted features or retraining a new network for each new subject—similar to best practices in face recognition systems [52, 40, 42]. (5) All studies use only one session for verification, even when multiple are available. They typically repeat experiments with different session pairs (one enrollment - one verification), averaging the results. In contrast, using the first session for enrollment and the rest for verification reflects a more realistic deployment setting.
@@ IV-E1 Comparison with Brainwave Authentications Systems
TABLE X: Performance metrics for different scenarios with detailed enrollment and verification session information. (Verification samples refer to the number of samples from the same session used in a single verification attempt.)
@@
First All Others 1 10.28±6.3010.28$\pm$ 6.3010.28 ± 6.30 43.33±24.4643.33$\pm$ 24.4643.33 ± 24.46 66.98±25.3066.98$\pm$ 25.3066.98 ± 25.30 80.65±21.4580.65$\pm$ 21.4580.65 ± 21.45
First All Others 4 6.22±5.646.22$\pm$ 5.646.22 ± 5.64 27.86±25.7527.86$\pm$ 25.7527.86 ± 25.75 50.39±30.4050.39$\pm$ 30.4050.39 ± 30.40 64.75±30.0364.75$\pm$ 30.0364.75 ± 30.03
First Second 1 6.87±7.616.87$\pm$ 7.616.87 ± 7.61 32.15±32.2932.15$\pm$ 32.2932.15 ± 32.29 54.04±35.6654.04$\pm$ 35.6654.04 ± 35.66 65.07±34.5265.07$\pm$ 34.5265.07 ± 34.52
First Second 4 3.04±6.233.04$\pm$ 6.233.04 ± 6.23 20.71±33.9820.71$\pm$ 33.9820.71 ± 33.98 33.44±40.5833.44$\pm$ 40.5833.44 ± 40.58 41.26±42.1641.26$\pm$ 42.1641.26 ± 42.16
First Two Third 1 4.87±5.364.87$\pm$ 5.364.87 ± 5.36 23.98±28.2923.98$\pm$ 28.2923.98 ± 28.29 43.59±35.0443.59$\pm$ 35.0443.59 ± 35.04 56.39±35.1556.39$\pm$ 35.1556.39 ± 35.15
First Two Third 4 1.54±2.741.54$\pm$ 2.741.54 ± 2.74 12.24±26.9912.24$\pm$ 26.9912.24 ± 26.99 24.28±35.3324.28$\pm$ 35.3324.28 ± 35.33 32.11±38.7932.11$\pm$ 38.7932.11 ± 38.79
First Two Third 16 0.68±1.750.68$\pm$ 1.750.68 ± 1.75 9.60±26.679.60$\pm$ 26.679.60 ± 26.67 16.08±33.3716.08$\pm$ 33.3716.08 ± 33.37 19.99±36.3119.99$\pm$ 36.3119.99 ± 36.31
IV-E2 Comparison with Biometric Industry Standards
[…]
@@ IV-E2 Comparison with Biometric Industry Standards
While our experiments show that brainwave authentication does not yet meet international standards, our previous experiment on the impact of headsets on performance (Sec.IV-B4) revealed that the headset with a higher number of subjects yielded a lower error rate, and we also know that biometrics that meet these standards typically employ datasets with thousands of subjects[53]. Therefore, we conducted an experiment to explore the relationship between the number of subjects used for training the feature extractor and the error rate. The results in Figure 8 indicate a linear relationship between the logarithm of the number of subjects and the EER. It is obvious that the linear decline eventually slows down. The current trend suggests that with a dataset of a few thousand subjects (from 2,000 to 16,000, depending on the device), it may be possible to approach biometric performance standards. This is reasonable when compared to the data volumes typically used in modalities like face recognition.
@@
Figure 8: EER vs. number of training subjects. Points show experimental results; dashed line shows linear prediction.
V Limitations
@@
Our study is subject to certain limitations. First, our stimuli were limited to a general text-based memorization task. Ideally, multiple task types could be included to allow for performance comparisons and identification of the best-performing task. Second, we simulated consumer-grade devices by selecting the same number and channel locations as those found in such devices. However, we did not have a measure to directly compare the quality of sensors. Although sensor quality will improve, the main challenge will remain their number and placement for effective use. It is worth noting that all other multi-session studies in the literature [16, 13, 11, 14] also employ medical-grade data. Finally, we only had access to a single dataset including a notable number of sessions and subjects. However, best practices in state-of-the-art face recognition research recommend training on multiple large datasets and testing on an independent dataset to ensure a more robust evaluation [54]. To simulate this, we split our dataset into a subject-independent set for training feature extraction and evaluation. Moreover, we partitioned the dataset based on device, showing that cross-device evaluation significantly increases error (Tables III and IV.) Achieving robust performance requires large amounts of data to extract distinctive features and account for subject-specific, environmental, and hardware-related variations. For example, in the case of face recognition, Kim et al. [54] used 3 datasets comprising over 350,000 subjects and more than 14.9 million images to train a generalized model. Such large-scale datasets are not yet available in brainwave authentication.
VI Conclusion and Future Work
This study investigated factors influencing the performance of brainwave authentication over multiple sessions. Our findings highlight that the pairwise deep learning approach for feature extraction outperforms hand-crafted features. Interestingly, we observed that larger evaluation sets do not necessarily make authentication more difficult; instead, they reduce the uncertainty of the results. We acknowledge that the error rate will eventually increase as the number of individuals grows, since only a limited number of individual templates can be reliably distinguished within a finite embedding space. Expanding the embedding space could mitigate this limitation, though evaluating this would require datasets with substantially more users.
@@ VI Conclusion and Future Work
Authentication performance remained relatively stable during the first week, with error rates increasing only slightly in the second month. However, error rates could increase by up to 2.1 times after one year, suggesting that the enrollment dataset should be periodically updated to maintain performance. We also examined the impact of recording devices, emphasizing the importance of including data from different headsets in training the feature extractor. Additionally, we showed that increasing the number of enrollment sessions and verification time can help reduce the EER, though only up to a certain point. Our findings also demonstrate that brainwave authentication remains feasible even with a reduced number of EEG sensors. Finally, our results are competitive with the state-of-the-art in brainwave authentication, investigated with stronger assumptions in simpler scenarios.
@@
For future work, there is an urgent need for a public multi-session dataset with a few thousand subjects to ensure brainwave authentication research meets international standards. This dataset should be collected using consumer-grade devices in multiple sessions, and incorporating multiple devices in the same session would be valuable for studying the cross-headset authentication challenge. On the technical side, further research should focus on developing deep learning architectures that better capture the unique features of brainwave data, as well as effective signal preprocessing approaches to remove noise. Adopting an open-source approach would enable researchers to build on each other’s work, promoting reproducibility and collaboration. Further investigation of more advanced attacks on brainwave authentication is necessary.
Acknowledgment
@@
This work was funded by the Topic Engineering Secure Systems of the Helmholtz Association (HGF) and supported by KASTEL Security Research Labs, Karlsruhe, and Germany’s Excellence Strategy (EXC 2050/1 ‘CeTI’; ID 390696704).